3. Trading in foreign exchange Aa Aa What are spot rates and forward rates? Suppose you open the newspaper today and observe the following indirect exchange rate quotations for the British pound Forward Exchange Rates 60 DaysS 0.5299 Spot Exchange Rates 30 Days 90 Days British pound (pound dollar) 0.5267 0.5283 0.5315 The British pound is selling at a in the forward market. Suppose you make a E 600,000 sale to a British customer who has 60 days to pay you in cash. The customer will pay you in British pounds, but your company is based in the United States, so you are most concerned with the dollar value of the payment. If the customer pays you E 600,000 today, how much is that worth in dollars? O $911,334 O $968,293 O $1,196,126 O $1,139,168 Assume that the forward market is correct and the 60-day forward exchange rate quoted in the newspaper today (above) is the spot exchange rate 60 days from now. If the customer waits the full 60 days and pays you E600,000, how much have you lost (in dollar terms) due to exchange rate fluctuations? $6,879 O $5,503 $6,535 O $5,159
Answer #1

1) The British pound is sold at a discount in the forward market

2) Worth of Euro in terms of dollars today = 6,000,000/0.5267 = 1,139,168(option d)

3) Worth of Euro in terms of dollars 60 days from today = 6,000,000/0.5299 = 1,132,289.11
Amount lost = 1,132,289.11 - 1,139,168 = -6879(Option a)

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well